# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v3

@@ 3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands to precharge the open row and fetch the target row to the row buffer. Therefore, a user space application can infer whether a row has been accessed recently by measuring the time it takes to access the row. To demonstrate the latency gap between different row buffer states, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts.
We observe that a row buffer conflict takes 74 CPU cycles more than a hit, which is large enough to detect. So, a user space application can exploit the row buffer as a timing covert and side channel. Our results align with the observations of prior works [68, 77, 78].
@@ 3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory. However, these covert-channel attacks have limited throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests.
There are several ways to bypass the cache hierarchy by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by (remote) direct memory access, (R)DMA, engines, and (iv) non-temporal memory hints (e.g., movnti). However, no attack primitive provides both fast and reliable communication.
@@
We identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties. First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications. Second, eviction sets require many load requests proportional to the number of ways in the caches, which scales with modern designs. Therefore, the attacks that use them have limited throughput. While the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other design choices that are not transparent to the attacker. Third, DMA and R-DMA engines incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example attack using this primitive[91] showcases a covert-channel attack based on memory contention. However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s). Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach. Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches. In contrast, PiM operations provide fast and reliable direct memory access to user space applications, which can be exploited for high-throughput covert- and side-channel attacks.
@@ 3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a realistic system (described in §5.1), establishing a state-of-the-art covert channel [68].
@@
Attack Scenarios. We simulate two attacks. First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets. To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line, and then it fetches the target row into the row buffer using an additional memory request. The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache. Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads. Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
@@
For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time needed to propagate bits from the sender to the receiver. We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
@@
Impact of LLC Size on Communication Throughput. Fig. 2 shows the impact of the LLC size on (i) the throughput of baseline and direct memory access attacks and (ii) the latency of a cache line eviction. The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [92].
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive. With increasing LLC size, the cache access latency increases and, thus, amplifies the cost of eviction, which requires multiple cache accesses.
@@
Figure 2: Impact of the LLC size on the data leakage throughput of covert channels and the eviction latency.
@@
Impact of LLC Associativity on Communication Throughput. Fig. 3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks and (ii) the eviction latency of a cache line. The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core). Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache. We observe that the baseline attack’s communication throughput decreases significantly with an increasing number of LLC ways due to the higher cache eviction latency.
@@
Figure 3: Impact of the number of LLC ways on the data leakage throughput of covert channels and the eviction latency.
@@
We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity. However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
The adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications. Therefore, it is crucial to carefully consider the security implications of PiM architectures and develop robust security measures to protect against potential attacks.
@@ 4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
This section introduces IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by the emerging processing-in-memory paradigm. IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. In this work, we present two covert-channel and one side-channel attacks enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions.
@@ 4.1 PnM-based IMPACT Covert-Channel Attack
Baseline PnM Architecture. PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture of the host processor with PIM-enabled instructions (PEIs). PEIs can be executed either on the host processor or on the in-memory computation logic. We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
The baseline PEI architecture has two key components. First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU. In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, the PEI architecture has a PEI Management Unit (PMU) that monitors data locality in application regions and decides where to map each region (either to host-side or memory-side PCUs). This way, PEI architecture can execute regions with high locality on the host side to benefit from caches. PEI adds an ignore flag per locality monitor entry to ignore the first hit in the locality monitor [93]. This is because if a locality monitor entry is allocated for a PIM operation, it is often too aggressive to consider the operation as having a high locality on its first hit. This bit allows IMPACT-PnM to easily bypass the locality monitor even with a small virtual address range at its disposal.
Attack Overview. Before the attack, the sender and the receiver co-locate their data in the same DRAM banks. To achieve this, one process uses memory massaging techniques to place its data in the same bank as the other process in main memory, as used in many prior works [76, 68, 94, 95].
IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank. In this attack, we encode logic-1 as interference in the DRAM bank (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
@@
1 Sender:
2 message = [0: N-1]; //N banks and N-bit msg.
3 batch_size = M; //M-bits per batch where M<N
4 init_semaphore(0);
5 barrier(); // synchronize with receiver
@@ 1 Sender:
9 if message[bank] == 1:
10 pim_add(row[bank]);
11 else:
@@ 12 NOP();
14 memory_fence(); // signal receiver
15 increment_semaphore(); // to start probing
16
17 Receiver:
18 init_row = initialize_DRAM_rows_with_PEIs();
@@ 17 Receiver:
20 barrier(); // synchronize with sender
21 for every batch:
22 block_on_semaphore();
23 //if semaphore==0, wait
@@
29 if latency > threshold:
@@
Step 1. The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 18). The receiver either uses unused cache lines of the activated row or sets the ignore flag in the corresponding entry in order to bypass the locality monitor. Then, the sender and the receiver synchronize on a barrier (Lines 5 and 20), which ensures that they start the communication turn at the same time.
Step 2. The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted. To transmit a 0 ( 1), the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains in the row buffer  2). To transmit a 1 ( 3), the sender activates a different row to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4). At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed and then signals the receiver to start probing the DRAM banks (Line 15).
Step 3. The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25). Then, the receiver measures the time it takes to complete the request (Line 27). If the latency exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6). Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8). The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
Sender-Receiver Synchronization. The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack. To do so, the sender and the receiver use a semaphore. The value of the semaphore indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed. The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing). When the sender completes the transmission of a batch, they increment the semaphore to signal the receiver to start. Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks. In parallel, the sender can start transmitting the next batch that utilizes a different set of DRAM banks. This way, they can overlap the transmission latency and the probing latency. When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted. In our attack implementation (§5.1), we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.

### v4

@@ 3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands (i.e., precharge the bank and fetch the target row). Therefore, a userspace application can infer whether a row has been accessed recently by measuring the time it takes to access the row.
To demonstrate the latency difference between a row buffer hit and a row buffer conflict in a PiM-enabled system, we simulate such a system (described in detail in §5.1). We run a microbenchmark that 1) creates memory accesses that result in row buffer hits and conflicts, and 2) measures the memory access latency for each memory request. We observe that a row buffer conflict takes 74 CPU cycles more than a hit for a CPU operating at 2.6 GHz, which is large enough to detect, which aligns with the observations of prior works [98, 107, 108]. Based on this observation, we conclude that a userspace application can exploit the row buffer as a timing channel in a PiM-enabled system.
@@ 3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory. However, these attacks have limited throughput in traditional architectures because accessing main memory directly is difficult due to deep cache hierarchies.
There are four main attack primitives that enables attackers to bypass the cache hierarchy and directly access main memory: 1) using specialized instructions to flush a target cache line (e.g., clflush [121]) from the cache deterministically, 2) using cache eviction sets [122] to replace a target cache line in the cache (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), 3) using a (remote) direct memory access engine (i.e., (R)DMA engine) [123, 124], and 4) using non-temporal memory hints (e.g., movnti instruction [121]).
We analyze the attack primitives in terms of efficiency and effectiveness. To do so, we identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives and PiM operations in terms of these properties.
@@
First, specialized cache flush instructions have two main disadvantages that limit their effectiveness at directly accessing main memory: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location [121], and 2) in some modern systems, these instructions are privileged and cannot be used by userspace applications (e.g., ARM processors [125]) (not shown in the table). Second, to directly access main memory, cache eviction sets require many load requests proportionally to the number of ways in the caches, which increases with modern designs. Therefore, the attacks that use them have limited throughput. Even though the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other memory system design choices that are not transparent to the attacker. Third, DMA and RDMA engines [123, 124] incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example work using this primitive [126] showcases a covert-channel attack based on memory contention. However, memory contention-based timing differences is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s). Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound by the instruction set architecture. These instructions may prefetch data from the last level cache (prefetchnta), or can be cached in intermediate buffers (e.g., movnti instruction) [127]. Consequently, the ISA does not guarantee that non-temporal hints will consistently and directly access main memory. In contrast, PiM operations provide fast and reliable direct memory access to userspace applications, which can be exploited for high-throughput covert- and side-channel attacks, as we show in this paper.
@@ 3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a PiM-enabled system (described in §5.1), establishing a state-of-the-art row buffer-based covert channel [98].
3.3.1 Attack Scenarios
@@
We simulate two attacks. First, we evaluate a baseline main memory attack that relies on cache bypassing using cache eviction sets. To transmit messages through the row buffer, the sender and receiver processes 1) evict the corresponding cache line to directly access main memory and 2) fetch the target row into the row buffer using an additional memory request. The baseline attack evicts a cache line by issuing N memory requests to the target cache set in an N-way cache. The actual eviction latency in a real system can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems [128, 129] or due to address translation overheads [130, 131, 132, 133, 134, 135, 136, 137]. Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
@@
For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time to propagate bits from the sender to the receiver. We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
3.3.2 Impact of LLC Size on Communication Throughput
@@
Fig. 2 shows the impact of the LLC size on (i) the throughput of the baseline and direct memory access attacks on the primary y-axis (left) and (ii) the latency of a cache line eviction on the secondary y-axis (right). The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). We calculate the cache access latency with increasing LLC sizes using CACTI 6.0 [138].
@@
Figure 2: Impact of the LLC size on data leakage throughput of covert channels and LLC eviction latency.
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline cache eviction sets based attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive. With increasing LLC size, the cache access latency increases, amplifying the latency of eviction, which requires multiple cache accesses. Therefore, we conclude that with increasing LLC sizes, cache access and eviction latencies increase, hindering the baseline attack’s communication throughput, compared to the direct memory access attack, which is independent of the LLC size.
3.3.3 Impact of LLC Associativity on Communication Throughput
@@
Fig. 3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks on the primary y-axis (left) and (ii) the eviction latency of a cache line on the secondary y-axis (right). The x-axis shows the increasing number of LLC ways (from 2 to 128) with a fixed total LLC size (e.g., 16 MB). We observe that as the number of LLC ways increases, the number of memory requests required to evict a cache line increases, greatly reducing the baseline attack’s communication throughput.
@@
Figure 3: Impact of LLC associativity on the data leakage throughput of covert channels and the eviction latency.
@@
Based on these two experiments, we conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity. In contrast, a covert channel based on direct memory access can achieve high throughput independently of the cache configuration, as it does not require bypassing the cache hierarchy.
@@ 4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by emerging PiM architectures. IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. In this work, we present two covert-channel and one side-channel attacks enabled by PnM and PuM solutions.
@@ 4.1 PnM-based IMPACT Covert-Channel Attack
Baseline PnM Architecture. PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture with instructions that can be executed in compute units in the processor core or near memory based on a dynamic decision made considering the locality of the required memory accesses. We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [89]).
The baseline PEI architecture has two key components. First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU. In-memory PCUs are shared among all cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, there is a PEI Management Unit (PMU) that monitors data locality in application regions with a locality monitor, and decides where to map each region (either to host-side or memory-side PCUs). This way, the PEI architecture can execute application regions with high data locality on the host side to benefit from the cache hierarchy. In contrast, if an application region does not benefit from the cache hierarchy, it is executed on the memory-side PCU to benefit from the lower memory access latency.
Attack Overview. Before the attack, the sender and the receiver co-locate their data in the same set of DRAM banks. To achieve this, one process uses memory massaging techniques to place its data in a page located in the same DRAM bank as the other process, as used in many prior works (e.g., [106, 98, 139, 140]). To incentivize the PnM architecture to execute the instruction in a memory-side PCU and directly access DRAM, the attack processes access different cache lines in the allocated rows, and allocate and use multiple rows in each bank.
IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank. In this attack, we encode logic-1 as interference (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
@@
1 Receiver (Step 1):
2 init_row = initialize_DRAM_rows_with_PEIs();
3
4 Sender (Step 2):
5 message[0: N-1]; // N-bit message
6 batch_size = M; // M-bit batch, where M<N
7 for every batch:
@@ 4 Sender (Step 2):
10 if message[bank] == 1: // activate a row
11 pim_add(sender_row[bank]);
12 else: // do not interfere with the receiver
@@ 13 NOP();
15 // ensure all memory requests are served
16 // before starting the next batch
17 memory_fence();
18
19 Receiver (Step 3):
20 for every batch:
@@ 19 Receiver (Step 3):
27 if latency > THRESHOLD: // row buffer conflict
@@
Step 1. The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 2, 1). Initialization copies the receiver’s data to row buffers.
Step 2. The sender transmits a batch of the message by mapping each bit in the batch to a separate DRAM bank ( 2). Then, the sender checks the bit value (Line 10). If the bit is 1, the sender issues a PEI to activate a row (Line 11, 3) and thus, copies its content to the row buffer ( 4). Otherwise, the sender issues a NOP instruction (Lines 12-13, 5) and does not copy any new data to the row buffer ( 6).
Step 3. The receiver probes the DRAM banks by issuing PEIs to the initialized rows and measuring the PEI latencies (Lines 23-25, 7). If the latency of a PEI exceeds a predetermined row buffer hit threshold, the receiver detects interference and decodes the bit as a 1 (Lines 27-28, 8). Otherwise, the access is considered a row buffer hit, and the receiver decodes the bit as a 0 (Lines 29-30, 9). After the transmission of a batch of the message (i.e., M bits), the sender and the receiver execute a memory fence (Lines 17 and 32) to ensure that all the PEIs are executed before moving on to the next batch.
The receiver accesses the next cache line in the initialized row to ensure its PEI is executed near memory by bypassing the locality monitor. If the number of batches is greater than the number of cache lines in a row, it uses another row in the same bank to measure access latency. The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
Sender-Receiver Synchronization. The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack. In our proof-of-concept implementation, the sender and the receiver use a semaphore. The semaphore’s value indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed. The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing). When the sender completes the transmission of a batch, they increment the semaphore to signal the receiver to start. Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks. When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted.